OpenStudy (anonymous):

what is the answer to 16 plus 28 divided 2 minus 6 over 10 minus 4 times 2

OpenStudy (anonymous):

\[16+28 = 44 2-\( 6 \10-8) = 2-(6\2) = 2-3= -1 44/-1 = -44\]